# Some applications of canonical metrics to Landau-Ginzburg models

### v1

@@ 2.2 Second step: resolving the base locus
Let $\mu\!:T_{\nabla}\to\mathfrak{g}^{\vee}$ denote a moment map for the action of the group $K$ of Hamiltonian isometries of $(T_{\nabla},\omega_{\nabla})$, normalised by $\int\langle\mu,v\rangle\omega^{2}_{\nabla}=0$ for all $v\in\mathfrak{g}$. By a special case of [37], Theorem 1, if the balancing condition
@@
where $E_{i}\subset Z^{\prime}$ denote the exceptional divisors of $p$, admits an extremal metric $\omega_{Z^{\prime},\varepsilon}$. But since the reduced automorphism group of $Z^{\prime}$ is trivial (since $p\!:Z^{\prime}\to T_{\nabla}$ blows up a point in each irreducible component of the toric boundary of $T_{\nabla}$), if follows that this extremal metric is in fact the unique cscK metric $\omega_{Z^{\prime},\varepsilon}$ representing its Kähler class.
@@ 2.2.1 Global quotient
Then, the balancing condition $\sum^{3}_{i=1}\mu(g\cdot p_{i})=0$ holds iff the kernel of the matrix $[\varphi_{i}(g\cdot p_{j})]$ contains the vector $(1,1,1)^{T}$. We compute explicitly the components
$([\varphi_{i}(p_{j})](1,1,1)^{T})_{1}$
@@
$([\varphi_{i}(p_{j})](1,1,1)^{T})_{2}=\frac{3\left(|\lambda|^{4}-|\nu|^{2}\right)\left(2e^{2a_{0}/3}|\nu|^{2}\left(|\lambda|^{4}+|\nu|^{2}\right)+e^{4a_{0}/3}|\lambda|^{2}+|\lambda|^{2}|\nu|^{6}\right)}{\left(|\lambda|^{4}+|\nu|^{2}\right)\left(e^{2a_{0}/3}+|\lambda|^{2}|\nu|^{2}\right)\left(e^{2a_{0}/3}|\lambda|^{2}+|\nu|^{4}\right)}.$
@@ 2.2.2 Kempf-Ness Theorem and Hilbert-Mumford Criterion
By the Kempf-Ness Theorem, the vanishing moment map condition is equivalent to the polystability of the $G$-orbit of the base locus $\operatorname{Bs}(\mathfrak{d}_{\Delta})$, regarded as a cycle, definining a point in the relevant Chow variety. By the Hilbert-Mumford Criterion, the latter can be checked using one-parameter subgroups (1PS) of $G$. The Hilbert-Mumford weight for the cycle is simply the sum of punctual Hilbert-Mumford weights (for more details see e.g. [35]).
@@
The computation of limit cycles as $t\to 0$ and Chow weights in the (nontrivial) chambers can be summarised as follows (we follow the convention that stable points have negative Chow weights):
@@ 2.3 Third step: crepant resolution
The final step in the canonical construction consists in passing to a crepant resolution of $Z^{\prime}$. In dimension $2$, this is given simply by two successive blowups of the orbifold points. Each singular point is replaced by two $(-2)$ curves. The pencil $\mathfrak{d}_{\Delta}$ induces a regular morphism $f\!:Z\to\mathbb{P}^{1}$, which is a rational elliptic surface and so a log Calabi-Yau compactified toric LG model for $\mathbb{P}^{2}$ (see e.g. [29], Section 3 or [18], Proposition 4.8).
@@ 3 $X=T_{P_{6a}}=S_{6}$
The degree $6$ del Pezzo $X=S_{6}$ is the unique smooth toric Kähler-Einstein del Pezzo surface. It is isomorphic to the blowup of $\mathbb{P}^{2}$ in three points in general position. We can choose a toric presentation such that the corresponding reflexive (smooth) polygon is given by $\Delta=P_{6a}$ in [18], Fig. 2. This in also the unique smooth del Pezzo polygon such that the polar dual $\nabla=(P_{6a})^{\circ}$ is equivalent to $P_{6a}$ up to a $GL(2,\mathbb{Z})$ transformation, so we have $T_{\nabla}\cong S_{6}$.
According to [29], Section 2, Facts 1.-2., $T_{\nabla}$ is anticanonically embedded in a projective space $\mathbb{P}^{6}$ with homogeneous coordinates $x_{i}$, $i=0,\cdots,6$ correponding to the integral points $c_{i}$ of $\Delta$, and with homogeneous ideal cut out by equations
@@
induced by monoid relations
@@
A toric LG potential $f_{S_{6},H}$ for $T_{\Delta}=S_{6}$ is given by Givental’s construction [16]. As before, we follow the inductive presentation given in [29], Section 3 (this is convenient for later purposes). So, we start from the potential for $S_{7}$, the blowup of $\mathbb{P}^{2}$ in two distinct points, with divisor
@@
So, we see that the Laurent polynomial $f_{S_{6},H}$ is the restriction to the structure torus of $T_{\nabla}$ of the birational map $T_{\nabla}\dashrightarrow\mathbb{P}^{1}$ defined by the pencil
@@ 3.1 Kempf-Ness Theorem
Theorem 1.5 now follows from the (generalised) Arezzo-Pacard Theorem, as discussed in detail in Section 2.
@@ 4 $T_{X}=T_{P_{9}}$
Then $P_{9}$ is obtained from $P_{6}$ be adding the vertices
@@ 4.1 Anticanonical class
since $\pi^{*}c_{1}(\mathcal{O}_{\mathbb{P}^{2}}(1))$ is pulled back from the base. ∎
@@
$E^{\prime 2}_{i,1}=-2,\,E^{\prime 2}_{i,2}=-1,\,E_{i,1}.E_{i,2}=1.$
@@ 4.2 Deformation theory
$E_{1}\mapsto\sum^{3}_{j=1}E_{1,j},\,E_{2}\mapsto\sum^{3}_{j=2}E_{1,j},\,E_{3}\mapsto E_{1,3},$
$E_{4}\mapsto\sum^{3}_{j=1}E_{2,j},\,E_{5}\mapsto\sum^{3}_{j=2}E_{2,j},\,E_{6}\mapsto E_{2,3},$
$E_{7}\mapsto\sum^{3}_{j=1}E_{3,j},\,E_{8}\mapsto\sum^{3}_{j=2}E_{3,j},\,E_{9}\mapsto E_{3,3}.$
@@
For $0<\varepsilon_{3}\ll\varepsilon_{2}\ll\varepsilon_{1}\ll 1$, depending only on $\delta>0$, such that the Kähler class
$\pi^{*}(\mathcal{O}_{\mathbb{P}^{2}}(1))-\varepsilon^{2}_{1}[E_{1}+E_{4}+E_{7}]-\varepsilon^{2}_{2}[E_{2}+E_{5}+E_{8}]-\varepsilon^{2}_{3}[E_{3}+E_{6}+E_{9}].$ (4.5)
@@
We need to show that for some $\delta>0$, sufficiently small, the open subset $\mathcal{D}_{>\delta}$ overlaps with an open neighbourhood of the class (4.4), consisting of classes which admit a unique cscK representative. This can be achieved by deforming the complex structure of $Y$ given by the iterated blowup.
@@
$H^{2}(Y,\mathcal{T}_{Y})\cong H^{2}(Y_{2},\mathcal{T}_{Y_{2}})\cong H^{2}(Y_{1},\mathcal{T}_{Y_{2}})\cong 0.$
So, small deformations of $Y$ correspond precisely to the blowup centers $\{p_{i}\}$, and the mirror map sends a neighbourhood of $c_{1}(S_{3})$ in the complexified Kähler cone holomorphically to a neighbourhood of $Y$ in its smooth Kuranishi space. On the other hand, the automorphism group of $Y$ is discrete. Thus, by the results of [15, 36], there is an open neighbourhood of $Y$ in its smooth Kuranishi space admitting a unique cscK metric in the class (4.5). For $\delta>0$ sufficiently small, the above neighbourhood of $c_{1}(S_{3})$ intersects $\mathcal{D}_{>\delta}$ in an open subset as required.
@@ 4.4 $K$-instability of adiabatic classes
for a fixed polarisation $H$ on $Y$ and with $k$ sufficiently large (depending on $(Y,H)$). These polarisations are very far from those considered in the present paper. Indeed, in the current case of the compactified toric Landau-Ginzburg $Y$ model mirror to a cubic surface $S_{3}$, it is even the case that the complex structure on $Y$ corresponding to $c_{1}(S_{3})\in\operatorname{Ka}_{\mathbb{C}}(S_{3})$ is $K$-unstable for all $k\gg 1$.
@@
Note that, for generic $[\omega_{\mathbb{C}}]\in\operatorname{Ka}_{\mathbb{C}}(S_{3})$, the corresponding fibration $f\!:Y\to\mathbb{P}^{1}$ has only fibres of type $I_{N}$ and so, by [22], Corollary 1.5 $(1)$, for any fixed polarisation $H$, the polarised surface $(Y,H+f^{*}c_{1}(\mathcal{O}_{\mathbb{P}^{1}}(k)))$ is $K$-stable for $k\gg 1$. The problem is that, by the above instability result, it is not possible to choose $H$ along a family of complex structures approaching the mirror to $c_{1}(S_{3})$ so that $k$ can be chosen uniformly. This problem does not arise for the polarisations we consider thanks to the uniform estimates in the Arezzo-Pacard and deformation theorems.
@@ 8 $T_{X}=T_{P_{8a}}$
are also cscK. By our deformation theory argument, for sufficiently small $\delta>0$, the generic set $\mathcal{D}_{>\delta}$ has nonempty intersection with an open neighbourhood of $c_{1}(S_{4})\in\operatorname{Ka}_{\mathbb{C}}(S_{4})$ corresponding to cscK complex structures on $Z$.
@@ 14 Quartic threefold $V_{4}$
The polar dual $\nabla=\Delta^{\circ}$ is indeed given by $P_{0}$, so $f$ extends to an anticanonical pencil on $T_{\nabla}\cong\mathbb{P}^{3}=\mathbb{P}[z_{0}:z_{1}:z_{2}:z_{3}]$. In this case it is easy to see this extension directly,
@@
Seyeddali and Székelyhidi [32] prove that the blowup of an extremal metric along a submanifold of codimension at least $2$ which is invariant under a maximal torus in the Hamiltonian isometry group is still extremal, in classes making the exceptional divisor sufficiently small. They also conjecture that the same result holds when the submanifold has codimension $2$. Thus, assuming their conjecture, $[\omega_{Z_{1}}]$ admits an extremal representative $\omega_{Z_{1}}$. On the other hand, according to [31], Section 5.2, the Futaki character of $Z_{1}$ vanishes identically on the Kähler cone of $Z_{1}$. Thus the extremal metric $\omega_{Z_{1}}$ is necessarily cscK.

### v2

@@ 2 $X=T_{P_{3}}=\mathbb{P}^{2}$
In this Section we start the case by case analysis which is necessary for the proof of Theorem 1.6, as outlined in Section 1.3. This is completed in Section 13.
@@ 2.2 Second step: resolving the base locus
Let $\mu\!:T_{\nabla}\to\mathfrak{k}^{\vee}$ denote a moment map for the action of the group $K$ of Hamiltonian isometries of $(T_{\nabla},\omega_{\nabla})$, with Lie algebra $\mathfrak{k}$, normalised by $\int\langle\mu,v\rangle\omega^{2}_{\nabla}=0$ for all $v\in\mathfrak{k}$.
Remark 2.1.
The extension of Kähler geometry to orbifolds (such as $T_{\nabla}$) is discussed in several references including [11], Section 1, [29], Section 3, [34], Section 1 and [38], Section 2.
By a special case of [42], Theorem 1, if the balancing condition
@@
where $E_{i}\subset Z^{\prime}$ denote the exceptional divisors of $p$, admits an extremal metric $\omega_{Z^{\prime},\varepsilon}$. But since the reduced automorphism group (i.e. the connected component of the identity in the group of Hamiltonian automorphisms) of $Z^{\prime}$ is trivial (since $p\!:Z^{\prime}\to T_{\nabla}$ blows up a point in each irreducible component of the toric boundary of $T_{\nabla}$, and there are no nontrivial holomorphic vector fields on $T_{\nabla}$ vanishing at all these blowup centres), it follows that this extremal metric is in fact the unique cscK metric $\omega_{Z^{\prime},\varepsilon}$ representing its Kähler class.
@@ 2.2.1 Global quotient
Then, we have
$\mu(g\cdot p_{j})=(\varphi_{1}(g\cdot p_{j}),\varphi_{2}(g\cdot p_{j}))^{T},$
so the balancing condition $\sum^{3}_{j=1}\mu(g\cdot p_{j})=0$ holds iff the kernel of the matrix $[\varphi_{i}(g\cdot p_{j})]$ contains the vector $(1,1,1)^{T}$. We compute explicitly the components
$([\varphi_{i}(g\cdot p_{j})](1,1,1)^{T})_{1}$
@@
$([\varphi_{i}(g\cdot p_{j})](1,1,1)^{T})_{2}=\frac{3\left(|\lambda|^{4}-|\nu|^{2}\right)\left(2e^{2a_{0}/3}|\nu|^{2}\left(|\lambda|^{4}+|\nu|^{2}\right)+e^{4a_{0}/3}|\lambda|^{2}+|\lambda|^{2}|\nu|^{6}\right)}{\left(|\lambda|^{4}+|\nu|^{2}\right)\left(e^{2a_{0}/3}+|\lambda|^{2}|\nu|^{2}\right)\left(e^{2a_{0}/3}|\lambda|^{2}+|\nu|^{4}\right)}.$
@@ 2.2.2 Kempf-Ness Theorem and Hilbert-Mumford Criterion
By the Kempf-Ness Theorem, the vanishing moment map condition is equivalent to the polystability of the $G$-orbit of the base locus $\operatorname{Bs}(\mathfrak{d}_{\Delta})$, regarded as a cycle, definining a point in the relevant Chow variety (see [31], Theorem 8.3 for a statement of the Kempf-Ness Theorem for arbitrary Kähler forms representing a rational polarisation). By the Hilbert-Mumford Criterion, the latter can be checked using one-parameter subgroups (1PS) of $G$. The Hilbert-Mumford weight for the cycle is simply the sum of punctual Hilbert-Mumford weights (for more details see e.g. [40]).
@@
The computation of limit cycles as $t\to 0$ and Chow weights in the (nontrivial) chambers can be summarised as follows (we use the convention that stable points have negative Chow weights):
@@ 2.3 Third step: crepant resolution
The final step in the canonical construction consists in passing to a crepant resolution of $Z^{\prime}$. In dimension $2$, this is given simply by blowing up the orbifold points twice. Each singular point is replaced by two $(-2)$ curves. The pencil $\mathfrak{d}_{\Delta}$ induces a regular morphism $f\!:Z\to\mathbb{P}^{1}$, which is a rational elliptic surface and so a log Calabi-Yau compactified toric LG model for $\mathbb{P}^{2}$ (see e.g. [33], Section 3 or [19], Proposition 4.8).
@@ 3 $X=T_{P_{6a}}=S_{6}$
The degree $6$ del Pezzo $X=S_{6}$ is the unique smooth toric Kähler-Einstein del Pezzo surface. It is isomorphic to the blowup of $\mathbb{P}^{2}$ in three points in general position. We denote the exceptional divisors by $E_{i}$. We can choose a toric presentation such that the corresponding reflexive (smooth) polygon is given by $\Delta=P_{6a}$ in [19], Fig. 2. This in also the unique smooth del Pezzo polygon such that the polar dual $\nabla=(P_{6a})^{\circ}$ is equivalent to $P_{6a}$ up to a $GL(2,\mathbb{Z})$ transformation, so we have $T_{\nabla}\cong S_{6}$.
According to [33], Section 2, Facts 1.-2., $T_{\nabla}$ is anticanonically embedded in a projective space $\mathbb{P}^{6}$ with homogeneous coordinates $x_{i}$, $i=0,\cdots,6$ corresponding to the integral points $c_{i}$ of $\Delta$, and with homogeneous ideal cut out by the equations
@@
induced by the monoid relations
@@
A toric LG potential $f_{S_{6},H}$ for $T_{\Delta}=S_{6}$ is given by Givental’s construction [17]. We follow the inductive presentation given in [33], Section 3 (this is convenient for later purposes). So, we start from the potential for $S_{7}$, the blowup of $\mathbb{P}^{2}$ in two distinct points, with divisor
@@
We see that the Laurent polynomial $f_{S_{6},H}$ is the restriction to the structure torus of $T_{\nabla}$ of the birational map $T_{\nabla}\dashrightarrow\mathbb{P}^{1}$ defined by the pencil
@@ 3.1 Kempf-Ness Theorem
The claim of Theorem 1.6 for our $X$ now follows from the (generalised) Arezzo-Pacard Theorem, as discussed in detail in Section 2.
@@ 4 $T_{X}=T_{P_{9}}$
Then $P_{9}$ is obtained from $P_{6}$ by adding the vertices
@@ 4.1 Anticanonical class
since $\pi^{*}c_{1}(\mathcal{O}_{\mathbb{P}^{2}}(1))$ is pulled back from the base, so the Chow weight of the cycle $\sum_{i}q^{\prime}_{i}$ with respect to $\pi^{*}c_{1}(\mathcal{O}_{\mathbb{P}^{2}}(1))$ equals the Chow weight of its projection $\sum_{i}q_{i}$ with respect to $c_{1}(\mathcal{O}_{\mathbb{P}^{2}}(1))$, and the latter Chow weight vanishes by polystability. ∎
@@
$E^{\prime 2}_{i,1}=-2,\,E^{\prime 2}_{i,2}=-1,\,E^{\prime}_{i,1}.E^{\prime}_{i,2}=1.$
@@ 4.2 Deformation theory
$E_{1}\mapsto\sum^{3}_{j=1}E_{1,j},\quad E_{2}\mapsto\sum^{3}_{j=2}E_{1,j},\quad E_{3}\mapsto E_{1,3},$
$E_{4}\mapsto\sum^{3}_{j=1}E_{2,j},\quad E_{5}\mapsto\sum^{3}_{j=2}E_{2,j},\quad E_{6}\mapsto E_{2,3},$
$E_{7}\mapsto\sum^{3}_{j=1}E_{3,j},\quad E_{8}\mapsto\sum^{3}_{j=2}E_{3,j},\quad E_{9}\mapsto E_{3,3}.$
@@
For $0<\varepsilon_{3}\ll\varepsilon_{2}\ll\varepsilon_{1}\ll 1$, depending only on $\delta>0$, the Kähler class
$\pi^{*}(\mathcal{O}_{\mathbb{P}^{2}}(1))-\varepsilon^{2}_{1}[E_{1}+E_{4}+E_{7}]-\varepsilon^{2}_{2}[E_{2}+E_{5}+E_{8}]-\varepsilon^{2}_{3}[E_{3}+E_{6}+E_{9}]$ (4.5)
@@
Proof.
We work in $\mathcal{D}_{>\delta}$ for a fixed $\delta>0$. Firstly, as we already observed, the Kähler class on $\operatorname{Bl}_{p_{1},p_{4},p_{7}}\mathbb{P}^{2}$
$\pi^{*}(\mathcal{O}_{\mathbb{P}^{2}}(1))-\varepsilon^{2}_{1}[E_{1}+E_{4}+E_{7}]$
is cscK for $0<\varepsilon_{1}\ll 1$. Let us denote by $K$ the compact torus fixing $p_{1}$, $p_{4}$, $p_{7}$ on $\mathbb{P}^{2}$. A straightforward application of the Hilbert-Mumford criterion shows that the cycle $p_{2}+p_{5}+p_{8}$ on $\mathbb{P}^{2}$ is polystable with respect to the full automorphism group of $(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(1))$ and so by continuity its Chow weight as a cycle in
$(\operatorname{Bl}_{p_{1},p_{4},p_{7}}\mathbb{P}^{2},\pi^{*}(\mathcal{O}_{\mathbb{P}^{2}}(1))-\varepsilon^{2}_{1}[E_{1}+E_{4}+E_{7}])$
with respect to any nontrivial 1PS in $K$ is strictly positive for a uniform $0<\varepsilon_{1}\ll 1$. Then, by the Arezzo-Pacard Theorem, the Kähler class on $\operatorname{Bl}_{p_{2},p_{5},p_{8}}\operatorname{Bl}_{p_{1},p_{4},p_{7}}\mathbb{P}^{2}$
$\pi^{*}(\mathcal{O}_{\mathbb{P}^{2}}(1))-\varepsilon^{2}_{1}[E_{1}+E_{4}+E_{7}]-\varepsilon^{2}_{2}[E_{2}+E_{5}+E_{8}]$
is cscK for all $0<\varepsilon_{2}\ll\varepsilon_{1}\ll 1$ depending only on $\delta>0$. The claim now follows since $\operatorname{Bl}_{p_{2},p_{5},p_{8}}\operatorname{Bl}_{p_{1},p_{4},p_{7}}\mathbb{P}^{2}$ has discrete automorphisms. ∎
Finally, we need to show that for some $\delta>0$, sufficiently small, the open subset $\mathcal{D}_{>\delta}$ overlaps with an open neighbourhood of the class (4.4), consisting of classes which admit a unique cscK representative. This can be achieved by deforming the complex structure of $Y$ given by the iterated blowup.
@@
$H^{2}(Y,\mathcal{T}_{Y})\cong H^{2}(Y_{2},\mathcal{T}_{Y_{2}})\cong H^{2}(Y_{1},\mathcal{T}_{Y_{2}})\cong 0$
(this follows e.g. from [21], Exercise 10.5 (b)). So, small deformations of $Y$ correspond precisely to the blowup centers $\{p_{i}\}$, and the mirror map sends a neighbourhood of $c_{1}(S_{3})$ in the complexified Kähler cone holomorphically to a neighbourhood of $Y$ in its smooth Kuranishi space. On the other hand, the automorphism group of $Y$ is discrete. Thus, by the results of [16, 28, 41], there is an open neighbourhood of $Y$ in its smooth Kuranishi space admitting a unique cscK metric in the class (4.5). For $\delta>0$ sufficiently small, the above neighbourhood of $c_{1}(S_{3})$ intersects $\mathcal{D}_{>\delta}$ in an open subset as required.
@@ 4.4 $K$-instability of adiabatic classes
This Section contains the proof of the Lemma stated in Remark 1.9.
@@
for a fixed polarisation $H$ on $Y$ and with $k$ sufficiently large (depending on $(Y,H)$). These polarisations are very far from those considered in the present paper. Indeed, in the current case of the compactified toric Landau-Ginzburg $Y$ model mirror to a cubic surface $S_{3}$, it is even the case that the complex structure on $Y$ corresponding to $c_{1}(S_{3})\in\operatorname{Ka}(S_{3})_{\mathbb{C}}$ is $K$-unstable for all $k\gg 1$.
@@
Note that, for generic $[\omega_{\mathbb{C}}]\in\operatorname{Ka}(S_{3})_{\mathbb{C}}$, the corresponding fibration $f\!:Y\to\mathbb{P}^{1}$ has only fibres of type $I_{N}$ and so, by [23], Corollary 1.5 $(1)$, for any fixed polarisation $H$, the polarised surface $(Y,H+f^{*}c_{1}(\mathcal{O}_{\mathbb{P}^{1}}(k)))$ is $K$-stable for $k\gg 1$. The problem is that, by the above instability result, it is not possible to choose $H$ along a family of complex structures approaching the mirror to $c_{1}(S_{3})$ so that $k$ can be chosen uniformly. This problem does not arise for the polarisations we consider thanks to the uniform estimates in the Arezzo-Pacard and deformation theorems.
@@ 8 $T_{X}=T_{P_{8a}}$
are also cscK. By our deformation theory argument, for sufficiently small $\delta>0$, the generic set $\mathcal{D}_{>\delta}$ has nonempty intersection with an open neighbourhood of $c_{1}(S_{4})\in\operatorname{Ka}(S_{4})_{\mathbb{C}}$ corresponding to cscK complex structures on $Z$.
@@ 14 Quartic threefold $V_{4}$
In this Section we begin the case by case analysis which leads to the proof of Theorem 1.11, following the outline given in Section 1.5. This is completed in Section 21.
@@
The polar dual $\nabla=\Delta^{\circ}$ is indeed given by $P_{0}$, so $f$ extends to an anticanonical pencil on $T_{\nabla}\cong\mathbb{P}^{3}=\mathbb{P}[z_{0}:z_{1}:z_{2}:z_{3}]$. In this case, it is easy to see this extension directly,
@@
Seyeddali and Székelyhidi [37] prove that the blowup of an extremal metric along a submanifold of codimension greater than $2$ which is invariant under a maximal torus in the Hamiltonian isometry group is still extremal, in classes making the exceptional divisor sufficiently small. They also conjecture that the same result holds when the submanifold has codimension $2$. Thus, assuming their conjecture, $[\omega_{Z_{1}}]$ admits an extremal representative $\omega_{Z_{1}}$. On the other hand, according to [36], Section 5.2, the Futaki character of $Z_{1}$ vanishes identically on the Kähler cone of $Z_{1}$. Thus the extremal metric $\omega_{Z_{1}}$ is necessarily cscK.